### v1

# Atomic Localization Fluorescent Microscopy

Abstract
Super-resolution microscopy has revolutionized the imaging of complex physical and biological systems by surpassing the Abbe diffraction limit. Recent advancements, particular in single-molecular localization microscopy (SMLM), have pushed localization below nanometer precision [1, 2, 3, 4, 5], by applying prior knowledge of correlated fluorescence emission from single emitters [6]. However, achieving a refinement from $1$ nm to $1$ngström demands a hundred-fold increase in collected photon signal [7, 6, 8]. This quadratic resource scaling imposes a fundamental barrier in SMLM, where the intense photon collection is challenged by photo-bleaching, prolonged integration times, and inherent practical constraints [9]. Here, we break this limit by harnessing the periodic nature of the atomic lattice structure. Specifically, applying this discrete grid imaging technique (DIGIT) in a quantum emitter system, we observe an exponential collapse of localization uncertainty once surpassing the host crystal’s atomic lattice constant. We further applied DIGIT to a large-scale quantum emitter array, enabling parallel positioning of each emitter through wide-field imaging. These results showcase that DIGIT unlocks a potential avenue to applications ranging from identifying solid-state quantum memories in crystals to the direct observation of optical transitions in the electronic structure of molecules.
preprint: APS/123-QED
Figure 1: DIGIT concept. a, Farfield PSF of an emitter with a diffraction-limited width $\sigma_{0}$ and its reconstructed location $(\mu_{x},\mu_{y})$ using SMLM. Multiple localizations’ distribution in one dimension in b, has a linewidth of $\sigma$. c, a zoomed-in view of the emitters’ underlying atomic structure - characterized by lattice constant a - illustrating their angström-level localization. d, transition of emitter localization from SMLM (purple) to DIGIT (green). e, DIGIT further extends in two dimensions. The color coding in SMLM represents distinct lattice groupings. f, The standard deviation of estimated location $\delta$: conventional SMLM in purple follows a polynomial scaling while DIGIT (green) scales exponentially as $\sigma$ falls below the lattice constant $a$
I Introduction
Optical super-resolution microscopy resolves the position of fluorescent emitters with a precision $\sigma$, that scales with the number of collected photons $N$. Specifically, single-molecular localization microscopy (SMLM), which determines emitters’ positions by reconstructing their far-field point-spread functions (PSFs) [6], is fundamentally limited by $\sigma\geq\sigma_{0}/\sqrt{N}$, where $\sigma_{0}$ is the standard deviation of the PSF, as shown in Fig. 1a. Recent works have achieved sub-nanometer $\sigma$ localizing both fluorescent dyes [2, 4] and quantum emitters [10, 5, 11]. However, these approaches assume that independent measurements are made of a continuous position variable $\mu_{x}$ (Fig. 1b). In many physical systems, emitters reside at discrete lattice sites, prompting a breakdown of the continuous-space assumption. For example, crystal defects such as color centers are confined to specific atomic positions in the host lattice (Fig. 1c) [12, 13]. As localization approaches the lattice scale, it becomes essential to incorporate these positional constraints. In this work, we introduce a super-resolution technique that explicitly factors in the discrete nature of emitter positions, resulting in a new localization scaling of $\sigma\propto e^{-\sqrt{N}}$. This exponential enhancement pushes beyond the conventional limit, enabling localization approaching the atomic regime.
To illustrate this concept, Fig. 1d presents the transition from continuous to discrete localization. The purple dashed line represents the conventional SMLM localization distribution $f(\mu_{x})$. By quantitatively linking the emitters’ underlying atomic lattice structure (black lines), the posterior localization distribution $P(x)\propto f(\mu_{x})\sum\delta(x-na)$ becomes discretized, sharply peaking at each lattice site $n$ with a periodicity $a$ (green lines). A detailed explanation of this incorporation of lattice constraints within a Bayesian framework is provided in Supplementary Information Sec. I.
The approach can be generalized to two dimensions where the position of each emitter is expressed not as $\bm{\mu}_{i}\pm\sigma_{i}$ but as an integer combination of lattice vectors: $\bm{\mu}_{i}=m_{i}\mathbf{a}_{1}+n_{i}\mathbf{a}_{2}$, where $(\mathbf{a}_{1},\mathbf{a}_{2})$ are lattice unit vectors and $(m_{j},n_{j})\in\mathbb{Z}$. We call this implementation the discrete grid imaging technique (DIGIT), and Fig 1e illustrates the resulting discrete localizations. The key advantage of DIGIT, as shown in Fig 1f, is that after $\sigma$ approaches the lattice constant $a$, the standard deviation of emitters’ estimated localization $\delta$ diminishes exponentially, beyond the polynomial scaling in conventional SMLM.
In this paper, we first demonstrate lattice-scale localization precision and accuracy using a diamond sample by reengineering the SMLM system. We then extend it to achieve atomic-scale localization using the Bayesian-based DIGIT approach. Additionally, we apply DIGIT to $>600$ emitters and examine the enhancement requirements in a large-scale quantum system. Collectively, these advancements establish DIGIT as a competitive tool for attaining unprecedented precise measurements, ultimately paving the way to direct optical resolution of crystal and atomic features within quantum systems.
Figure 2: Sub-diffraction imaging of SiVs. a, the schematic of sample preparation and characterization, including SiV FIB implantation and PL collection at cryogenic temperatures. b, Diffraction-limited PL image of implanted SiV clusters. c, PLE spectrum of the cluster highlighted in b, with Lorentzian fitting shown in green. The top inset displays an EMCCD image of resolved emitter 1 with scale bar of $200$ nm. d,e, Comparison of SMLM-reconstructed images without and with digital twin calibration. In d, closely spaced emitters appear unresolved, whereas in e, distinct emitters emerge with a localization uncertainty towards $\sigma=0.6$. f, The localization precision $\sigma$ improves following shot noise limited scaling.
II Lattice-scale Accuracy and Precision
Diamond color centers (CCs) serve as exemplary systems for DIGIT because of their long-lived emission and well-defined locations within the diamond crystal lattice. Although previous studies have achieved sub-nanometer localization $\sigma$ in diamond CCs [14, 15, 5, 1], several challenges remain in their imaging accuracy. These include mitigating the biases of the fitting model with fixed dipole nature of atomic transitions [16, 17, 18, 19] and drift-induced localization errors [20]. To address these challenges, we co-designed an end-to-end physical-digital twin system for SMLM, which accurately digitizes the experimental apparatus and optimizes its performance (Extended Data Fig. 2). By integrating this modeling into the microscope design and data acquisition, we achieve lattice-scale accuracy and precision.
The physical system involves a custom widefield cryogenic photoluminescence (PL) microscope (Fig. 2a) to image an array of silicon-vacancy centers (SiV) produced by focused-ion beam (FIB) implantation (See Methods). The farfield emission from SiVs is collected by an electron-multiplying charge-coupled device (EMCCD) camera. Fig. 2b shows the resulting PL image of SiV cluster arrays from off-resonant excitation.
To resolve individual SiV emitters beyond the diffraction limit, we applied prior knowledge of the natural inhomogeneous broadening of SiV transition energies [21] as illustrated in Fig. 2a. Specifically, we employed cryogenic photoluminescence excitation (PLE) by sweeping the excitation laser frequency across the zero-phonon line transitions and collecting light emitted into the phonon sideband on the EMCCD camera (see Methods). Fig. 2c presents the PLE spectrum of the cluster highlighted in Fig. 2b, revealing three distinct peaks with extracted resonance frequencies at $\omega_{1}/2\pi=406.708$ THz, $\omega_{2}/2\pi=406.724$ THz, $\omega_{3}/2\pi=406.735$ THz. With the laser frequency locked at each $\omega_{i}$, we subsequently re-acquired EMCCD images at extended integration times. For example, Fig. 2c shows an EMCCD image of SiV 1’s PSF with $\omega_{1}$.
Next, we developed a comprehensive digital twin model to extract CCs’ positions from their emission profiles. Instead of a simplified 2D Gaussian fit, our digital twin accurately represents the entirety of the physical apparatus, correcting for system-induced biases to achieve Ångström-level localization accuracy and precision. Built on a fixed-dipole model governing far-field dipole emissions [22, 23], it further incorporates imperfections of the PLE setup such as optical aberrations, interpixel non-uniformity, electronic noise in the EMCCD, and setup drift. We verified the model’s accuracy by examining two orthogonally polarized optical transitions within a single SiV [21] and observed the two distinct PSFs overlap. Further details are provided in Supplementary Information Sec. III.
Fig. 2d and e compare the resulting localization probability density function $f(\mu_{i})$ derived from (d) a basic 2D Gaussian fit and (e) our digital twin model. In fig. 2d, the simple Gaussian approach yields a localization uncertainty of $\sigma=1.8$ nm, insufficient to resolve two closely spaced emitters. However, by applying the accurate modeling, $\sigma$ reduces to $0.6$ with integrating PSF over $8000$ frames, as illustrated in Fig. 2e. Importantly, the three CCs are clearly spatially resolved. Fig. 2f plots the uncertainty $\sigma$ of the highlighted emitter in Fig. 2e as a function of integration time $t$. The standard error follows the shot-noise-limited scaling [24]. As $t$ increases, $\sigma$ decreases past the threshold at the diamond lattice constant $a$. The localization results for three emitters are provided in Supplementary Information Sec. IV.
Figure 3: Atomic-scale precision achieved by DIGIT. a, Comparison of SMLM localization distribution $f(x)$ in purple versus DIGIT posterior distribution $P(x)$ in green, aligned along the lattice by grey dots. The inset illustrates the diamond lattice model projected along the [100] crystallographic direction. b, Standard deviation of the mean distance between all emitter pairs. The experimentally measured localization precision using DIGIT(green) with its simulation predictions(yellow), demonstrating an exponential improvement over the SMLM shot noise scaling.
III Atomic Localization with DIGIT
Figure 4: Large-scale DIGIT. a, Three-dimensional representation of widefield PLE obtained by sweeping a resonant laser over 30 GHz in a $10$ MHz increment. b, Reconstructed clusters, each color-coded by the number of emitters $M$. c, Histogram of measured precision $\sigma$ for individual resolved emitters whose $\sigma<0.3a$. d,e, DIGIT performance under varying cluster density conditions, demonstrated through numerical simulations (upper panels) and experimental results (lower panels). d studies DIGIT’s advantage over SMLM by evaluating $\delta_{P}/\delta_{f}$ at different emitter counts $M$ per cluster and e studies the effect of cluster size $L$ when $M=5$.
Building on the localization data from the previous section, we applied DIGIT to image SiVs with a Bayesian prior that encodes the diamond lattice coordinates. Since the PSF is collected along diamond’s [100] crystallographic direction, we simplified the intricate three-dimensional lattice by projecting it onto a two-dimensional plane along the [100] axis, as shown in Fig. 3a. We then computed $P(x)$ after fitting the lattice positions via an affine transformation whose parameters are optimized through maximum likelihood estimation (see Extended Data Fig. 3).
Fig. 3a illustrates the resulting lattice positions (grey dots) and the new SiV localization $P(x)$ determined by DIGIT. In contrast to the purple contour representing the conventional SMLM localization distribution $f(\mu)$, DIGIT localizations $P(x)$ (green dots) exhibit discrete peaks that align well with specific lattice sites, reflecting the Bayesian incorporation of diamond’s periodic lattice structure.
Fig. 3b quantifies DIGIT by looking at the average localization deviation $\delta(r)$ over three emitter pairs at different stages of $\sigma$. We compare the DIGIT results (green dots) with conventional SMLM represented (purple dashed line) and the Monte Carlo simulation of the DIGIT-based digital twin (yellow line). Our experimental result of $\delta(r)$ aligns closely with the DIGIT simulation, surpassing the conventional limit of SMLM. Error bars represent the confidence intervals of localization uncertainty. The inset of Fig. 3b zooms in on the regime where $\sigma<0.35a$ DIGIT enhancement becomes pronounced. Each green dot represents an individual measurement at a specific $\sigma$. As a result, DIGIT enables localization entering an atomic regime of $\delta_{P}(r)=0.178\pm 0.107$ below the diamond lattice spacing, while conventional SMLM yields a mean uncertainty of $\delta_{f}(r)=0.7$ for the same three emitter pairs. This 5-fold improvement over SMLM enables atomic resolution that would otherwise necessitate integration time 25 times longer.
IV Large-scale DIGIT
To explore the full capability of DIGIT, we extend the technique from one quantum cluster to multiple clusters across $20$x20202020 $\mu$m field of view. Using widefield PLE, we successfully resolved over $600$ single emitters within $200$ clusters. The localization of these emitters was facilitated by the pre-etched markers in diamond (Extended Data Fig. 1), which served as defined reference coordinates. Expanding to a wider field of view emphasizes the importance of the implementation of aberration corrections and camera interpixel non-uniformity calibrations to maintain high localization precision (see Supplementary Information Sec. III). Through the analysis of various clusters, we investigated the influence of parameters enabled by widefield measurements, particularly focusing on clusters’ density of states and brightness.
Fig. 4a presents a hyperspectral image of widefield PLE, spanning spatial $(x,y)$ and atomic transition frequency $\omega/2\pi$. We applied intensity-based region grouping to the PLE EMCCD image to distinguish individual emitters whose transitions range from $406.69$ to $406.73$ THz. Fig. 4b shows the reconstructed cluster density map, with each cluster color-coded by the emitter counts $M$ identified through PLE spectrum (See supplementary Information Sec. V). A detailed flowchart of the widefield PLE data analysis can be found in Extended Data Fig. 4.
Next we integrated $8000$ frames at each resolved frequency $\omega_{i}$ and measured the localization uncertainty $\sigma$ for each emitter, as summarized in Fig. 4c. Only CCs exhibiting sub-Ångström $\sigma$ were counted for subsequent analysis. Emitters that did not meet this requirement have typically reduced brightness, exhibit blinking behavior [21], or experience non-uniformity widefield illumination [25], all of which contribute to worse localization uncertainty.
Figs. 4d and 4e present both numerical and experimental studies of the localization error enhancement $\delta_{P}/\delta_{f}$ across varying cluster densities. In Fig. 4d, we evaluate DIGIT’s performance with different emitter counts $M$ within a cluster. The simulation predicts that $\delta_{p}/\delta_{f}=1/M$, indicating that the localization error can be effectively compensated for by higher emitter counts. This is because additional emitters help to refine lattice alignment through MLE. Experimentally, we examined three cluster densities $M=3,4,5$, each closely matching the simulation prediction. This result confirms the collective advantage of dense-implanted clusters.
In Fig. 4e, we investigate $\delta_{P}/\delta_{f}$ for clusters with varying spatial separations $L$ at $M=5$. Numerical simulations show that DIGIT’s advantage diminishes as L grows, because distinguishing the maximum likelihood lattice point from adjacent points requires increasingly precise estimates of the rotation angle $\theta$, as $L*\theta\leq a$. Achieving such high precision poses significant challenges for MLE methods due to computational and experimental limitations. Experimentally, this scenario is further complicated by strain-induced variations in lattice coherence length, which can result in uneven lattice structures and gaps that impede precise localization [26].
Nevertheless, by extending DIGIT to $>600$ emitters across multiple clusters, we demonstrate its scalability and efficiency for parallel analysis. In contrast to scanning-based approaches that image each emitter individually, DIGIT processes many emitters simultaneously, achieving high-throughput localization.
V Discussion and Outlook
As the localization precision of conventional super-resolution microscopy approaches the length scale of electronic orbitals coupled to optical fields, it becomes essential to incorporate the atomic structure of samples as a Bayesian prior. Our demonstration of the DIGIT was made possible by achieving localization precision ($\sigma$) comparable to the lattice constant ($a$), allowing emitters to be accurately mapped to discrete lattice points rather than being confined to a continuous spatial domain. Furthermore, DIGIT provides a new precision scaling law: $\delta_{P}\propto e^{-\sqrt{N}}$ that is isotropic in all dimensions. Besides, parallelized DIGIT demonstrates its scalability by simultaneously localizing multiple emitters. To further access denser clusters, improved spectral resolution becomes critical. One promising strategy involves operating PLE without a repump laser so that emitters approach lifetime-limited spontaneous emission [27, 28].
Looking ahead, DIGIT has a direct application in diamond-based quantum information processes. For instance, the ability to achieve sub-Ångström localization enables in-situ tracking of emitters’ diffusion and interaction [29, 30, 31, 32], increases spatial storage density [33], and facilitates the generation and control of large cluster states made of locally coupled spins [34, 35]. Additionally, it is readily transferable to other periodic host materials, aiding to identify solid-state quantum memories. Potential candidates include point-like defects [36, 37, 38]; extended electronic orbitals in semiconductor quantum dots [39, 40], or molecules such as dibenzoterrylene, tetracene, or other complexes [41, 42, 43, 44, 45].
However, DIGIT is currently constrained to applications with well-characterized lattice structures. To overcome this limitation, our theoretical proposal for diamond suggests that with higher density CC, the atomic structure and the associated electronic orbital transitions can be learned, all from optical observation with zero-knowledge base (see Supplementary Information Sec. VI). Advanced computational techniques, such as deep learning of PSF from experimental observations [46, 47] and surrogate models [48], or generative AI methods such as adaptive model-based measurements, promise further acceleration.
Acknowledgments –We thank M.Mazaheri, Q.Peng for valuable input with the digital twin localization model. We thank C.S. Peng, A. Roy, H. Wang, J. Ren, Y. Liu for helpful discussions. We thank M.Mazaheri and V.Saggio for proofreading the manuscript. We acknowledge excellent technical support at the Massachusetts Institute of Technology (MIT) for MIT.Nano. Y.Yu performed FIB implantation, J. Daley performed evaporation medal growth. This work was supported in part by the National Science Foundation (NSF) Science-Technology Center (STC) Center for Integrated Quantum Materials (CIQM) under Grant No. DMR-1231319, by the NSF Engineering Research Center for Quantum Networks (CQN) awarded under Cooperative Agreement No. 1941583, and by the MITRE Moonshot Program. Y.D. acknowledges support by Mathworks fellowship.
Author Contribution –Y.D. conducted diamond fabrication, SiV implantation, digital twin development, localization analysis and measurements of the whole experiment. K.C.C. assisted with diamond fabrication. M.E.T assisted with building the wide-field SMLM, the dipole radiation model and digital twin verification. Q.G, Y.H, and D.R.E assisted with calibrating the camera photoresponse. Y.D and Q.G developed and implemented the Bayesian framework. Y.D, M.E.T and D.R.E conducted the theory proposal of optically resolved lattice structure. Y.D and D.R.E. conceived and supported the project. Y.D., M.E.T. and D.R.E wrote the manuscript with input from all authors.
Competing interests –The authors declare no competing financial interests.
Data availability –All the data that support the findings of this study are included in the Article and its Supplementary Information. Source data are available via figshare at 10.6084/m9.figshare.28281212
Code availability – Code for DIGIT principle and widefield DIGIT is available at: https://github.com/sophiaOnPoint/DIGIT